OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sec^2 x-2=tan^2 x

OpenStudy (anonymous):

\[\sec^2(x)-2=\tan^2(x)\]Is this how it's supposed to be? @mangnagelytis

geerky42 (geerky42):

Remember, \(\tan^2 x + 1 = \sec^2 x\), so \(\sec^2 x - 2 = \tan x = \sec^2 - 1\) \[\sec^2 - 2 = \sec^2 - 1\] No solution.

geerky42 (geerky42):

Oops I mean sec²x − 2=sec²x − 1

OpenStudy (anonymous):

This is all based on the assumption that what I typed is what the actual problem is. @geerky42 And in that case, there is no solution. @mangnagelytis

OpenStudy (anonymous):

That's what I thought. Thank you!